OpenStudy (niksbhalla14mar):

progression series x=1/80(41) + 1/79(42) + 1/78(43)..... + 1/41(80) and y=1-1/2+1/3-1/4+1/5......+1/79 - 1/80 solve for y/x

OpenStudy (3mar):

y/x=ln2/43.1193=0.0161

OpenStudy (niksbhalla14mar):

firstly, whaaaaaaaaaaaaat ? How did you solve that ? Secondly, answer is wrong..

OpenStudy (3mar):

If you are not sure, calculate it manually, you will get the same result. I know that you won't try. Just in case

OpenStudy (kevin):

@3mar How did you get that?

OpenStudy (3mar):

To be honest, I did it manually. Do it yourself, you get what I got.

OpenStudy (kevin):

lol

OpenStudy (3mar):

Can you?

OpenStudy (kevin):

No, explain to me

OpenStudy (niksbhalla14mar):

Sorry dude... you wont get the answer as what you got... the answer is 60.5

OpenStudy (3mar):

Can you explain how did you get the answer, Mr niksbhalla14mar?

OpenStudy (niksbhalla14mar):

sure... if you say please...

OpenStudy (niksbhalla14mar):

@3mar

OpenStudy (niksbhalla14mar):

Please find the attacched solution @3mar

OpenStudy (3mar):

What about this?

OpenStudy (3mar):

y series

OpenStudy (niksbhalla14mar):

Please see, i have solved the same. This is from logic... not from derived works.. We needed to find y/x , not y alone.

OpenStudy (3mar):

You got y=121 and I got it ln2 who is the correct?

OpenStudy (sshayer):

y=ln 2 if terms are infinite.

OpenStudy (3mar):

So if it is not infinite, it becomes 121 toooo big is not it?

OpenStudy (sshayer):

\[1-\left( \frac{ 1 }{ 2 }-\frac{ 1 }{ 3 } \right)-\left( \frac{ 1 }{ 4 }-\frac{ 1 }{ 5 } \right)-...\left( \frac{ 1 }{ 78 }-\frac{ 1 }{ 79 } \right)-\frac{ 1 }{ 80 } <1\]

OpenStudy (raden):

x=1/(80*41) + 1/(79*42) + 1/(78*43) + ... + 1/(41*80) x = 2 [1/(80*41) + 1/(79*42) + 1/(78*43) + ... + 1/(61*60)] and y=1 - 1/2 + 1/3 - 1/4 + .... + 1/79 - 1/80 y = (1 + 1/2 + 1/3 + .... + 1/79 + 1/80) - 2(1/2 + 1/4 + 1/6 + ... + 1/80) y = (1 + 1/2 + 1/3 + .... + 1/79 + 1/80) - 2*1/2 (1 + 1/2 + 1/3 + ... + 1/40) y = (1 + 1/2 + 1/3 + .... + 1/79 + 1/80) - (1 + 1/2 + 1/3 + ... + 1/40) y = 1/41 + 1/42 + 1/43 + ... + 1/80 y = (1/41+1/80) + (1/42+1/79) + (1/43+1/78) + ... (1/60+1/61) y = 121/(41*80) + 121/(42*79) + 121/(43*78) + ... + 121/(60*61) y = 121 [1/(41*80) + 1/(42*79) + 1/(43*78) + ... + 1/(60*61)] Thus, y / x = 121 [1/(41*80) + 1/(42*79) + 1/(43*78) + ... + 1/(60*61)] divided by 2 [1/(80*41) + 1/(79*42) + 1/(78*43) + ... + 1/(61*60)] = 121/2